OpenStudy (anonymous):

The principal P is borrowed and the loan's future value, A, at time t is given. Determine the loan's simple interest rate, r, to the nearest tenth of a percent. P = $10,000, A = $14,060, t = 2 years

OpenStudy (anonymous):

So confused, please help!

OpenStudy (anonymous):

I got 12,180 but I am pretty sure that's wrong.

hero (hero):

A = P(1+rt) 14,060 = 10,000(1+2r) 14,060/10,000 = 1 + 2r ((14,060/10000)-1)/2 = r

OpenStudy (anonymous):

oh my gosh, thank you! I appreciate you taking a look at my problem!

OpenStudy (anonymous):

Can you take a look at another for me, please?

OpenStudy (anonymous):

I'll post it.